Solve for x: ln(x^2-12)= ln(x)
Answer by user_dude2008(1862)

ln(x^2-12)= ln(x)


x^2-12=x

x^2-x-12=0

(x-4)(x+3)=0

x-4=0 or x+3=0

x=4 or x=-3 (ignore)

Only answer is x=4